\begin{thebibliography}{10}

\bibitem{A14}
J.~Aberg.
\newblock Catalytic coherence.
\newblock {\em Phys. Rev. Lett.}, 113:150402, 2014.

\bibitem{BC14}
T.~Baumgratz, M.~Cramer, and M.~B. Plenio.
\newblock Quantifying coherence.
\newblock {\em Physical review letters}, 113(14):140401, 2014.

\bibitem{ChDAP08}
G.~Chiribella, G.~M. D'Ariano, and P.~Perinotti.
\newblock Transforming quantum operations: Quantum supermaps.
\newblock {\em Europhysics Letters}, 83(3):30004, 2008.

\bibitem{ChG16}
E.~Chitambar and G.~Gour.
\newblock Comparison of incoherent operations and measures of coherence.
\newblock {\em Physical Review A}, 94(5):052336, 2016.

\bibitem{ChG16-2}
E.~Chitambar and G.~Gour.
\newblock Critical examination of incoherent operations and a physically
  consistent resource theory of quantum coherence.
\newblock {\em Physical review letters}, 117(3):030401, 2016.

\bibitem{CH16}
E.~Chitambar and M.~H. Hsieh.
\newblock Relating the resource theories of entanglement and quantum coherence.
\newblock {\em Physical review letters}, 117(2):020402, 2016.

\bibitem{CMM18}
T.~Cooney, M.~Mosonyi, and M.~M. Wilde.
\newblock Strong converse exponents for a quantum channel discrimination
  problem and quantum-feedback-assisted communication.
\newblock {\em Communications in Mathematical Physics}, 344(3):797--829, 2016.

\bibitem{C15}
P.~Cwikliski, M.~Studziski, M.~Horodecki, and J.~Oppenheim.
\newblock Limitations on the evolution of quantum coherences: towards fully
  quantum second laws of thermodynamics.
\newblock {\em Physical review letters}, 115(21):210403, 2015.

\bibitem{DevS16}
J.~I. De~Vicente and A.~Streltsov.
\newblock Genuine quantum coherence.
\newblock {\em Journal of Physics A: Mathematical and Theoretical},
  50(4):045301, 2016.

\bibitem{DBG15}
S.~Du, Z.~Bai, and Y.~Guo.
\newblock Conditions for coherence transformations under incoherent operations.
\newblock {\em Physical Review A}, 91(5):052120, 2015.

\bibitem{G63}
R.~J. Glauber.
\newblock Coherent and incoherent states of the radiation field.
\newblock {\em Physical Review}, 131(6):2766, 1963.

\bibitem{GW21}
G.~Gour and M.~M. Wilde.
\newblock Entropy of a quantum channel.
\newblock {\em Physical Review Research}, 3(2):023096, 2021.

\bibitem{HH18}
M.~L. Hu, X.~Hu, J.~Wang, Y.~Peng, Y.~R. Zhang, and H.~Fan.
\newblock Quantum coherence and geometric quantum discord.
\newblock {\em Physics Reports}, 762:1--100, 2018.

\bibitem{LKDW18}
F.~Leditzky, E.~Kaur, N.~Datta, and M.~M. Wilde.
\newblock Approaches for approximate additivity of the holevo information of
  quantum channels.
\newblock {\em Physical Review A}, 97(1):012332, 2018.

\bibitem{LY20}
Y.~Liu and X.~Yuan.
\newblock Operational resource theory of quantum channels.
\newblock {\em Physical Review Research}, 2(1):012035, 2020.

\bibitem{L15}
M.~Lostaglio and T.~Jennings, D.~Rudolph.
\newblock Description of quantum coherence in thermodynamic processes requires
  constraints beyond free energy.
\newblock {\em Nat. Commun.}, 6:6383, 2015.

\bibitem{MY16}
J.~Ma, B.~Yadin, D.~Girolami, V.~Vedral, and M.~Gu.
\newblock Converting coherence to quantum correlations.
\newblock {\em Physical review letters}, 116(16):160407, 2016.

\bibitem{RPJ16}
C.~Radhakrishnan, M.~Parthasarathy, S.~Jambulingam, and T.~Byrnes.
\newblock Distribution of quantum coherence in multipartite systems.
\newblock {\em Physical review letters}, 116(15):150504, 2016.

\bibitem{RPL16}
S.~Rana, P.~Parashar, and M.~Lewenstein.
\newblock Trace-distance measure of coherence.
\newblock {\em Physical Review A}, 93(1):012110, 2016.

\bibitem{R16}
A.~E. Rastegin.
\newblock Quantum-coherence quantifiers based on the tsallis relative $\alpha$
  entropies.
\newblock {\em Physical Review A}, 93(3):032136, 2016.

\bibitem{RM09}
P.~Rebentrost, M.~Mohseni, and A.~Aspuru-Guzik.
\newblock Role of quantum coherence and environmental fluctuations in
  chromophoric energy transport.
\newblock {\em J. Phys. Chem. B}, 113:9942, 2009.

\bibitem{SCG20}
G.~Saxena, E.~Chitambar, and G.~Gour.
\newblock Dynamical resource theory of quantum coherence.
\newblock {\em Physical Review Research}, 2(2):023298, 2020.

\bibitem{SZ97}
M.~O. Scully and M.~S. Zubairy.
\newblock Quantum optics.
\newblock {\em (Cambridge.) Ch}, 4:17, 1997.

\bibitem{SX15}
L.~H. Shao, Z.~Xi, H.~Fan, and Y.~Li.
\newblock Fidelity and trace-norm distances for quantifying coherence.
\newblock {\em Physical Review A}, 91(4):042120, 2015.

\bibitem{SAP17}
A.~Streltsov, G.~Adesso, and M.~B. Plenio.
\newblock Colloquium: Quantum coherence as a resource.
\newblock {\em Reviews of Modern Physics}, 89(4):041003, 2017.

\bibitem{SS15}
A.~Streltsov, U.~Singh, H.~S. Dhar, M.~N. Bera, and G.~Adesso.
\newblock Measuring quantum coherence with entanglement.
\newblock {\em Physical review letters}, 115(2):020403, 2015.

\bibitem{TK16}
K.~C. Tan, H.~Kwon, C.~Y. Park, and H.~Jeong.
\newblock Unified view of quantum correlations and quantum coherence.
\newblock {\em Physical Review A}, 94(2):022329, 2016.

\bibitem{TEZP19}
T.~Theurer, D.~Egloff, L.~Zhang, and M.~B. Plenio.
\newblock Quantifying operations with an application to coherence.
\newblock {\em Physical review letters}, 122(19):190405, 2019.

\bibitem{V22}
A.~Vershynina.
\newblock Measure of genuine coherence based of quasi-relative entropy.
\newblock {\em Quantum Information Processing}, 21(5):1--22, 2022.

\bibitem{V23}
A.~Vershynina.
\newblock Coherence as entropy increment for tsallis and r{\'e}nyi entropies.
\newblock {\em Quantum Information Processing}, 22(2):127, 2023.

\bibitem{Wat18}
J.~Watrous.
\newblock {\em The theory of quantum information}.
\newblock university press, Cambridge, 2018.

\bibitem{WWY14}
M.~M. Wilde, A.~Winter, and D.~Yang.
\newblock Strong converse for the classical capacity of entanglement-breaking
  and hadamard channels via a sandwiched r{\'e}nyi relative entropy.
\newblock {\em Communications in Mathematical Physics}, 331:593--622, 2014.

\bibitem{WY16}
A.~Winter and D.~Yang.
\newblock Operational resource theory of coherence.
\newblock {\em Physical review letters}, 116(12):120404, 2016.

\bibitem{WM13}
B.~Witt and F.~Mintert.
\newblock Stationary quantum coherence and transport in disordered networks.
\newblock {\em New J. Phys.}, 15:093020, 2013.

\bibitem{YZ16}
X.~D. Yu, D.~J. Zhang, G.~F. Xu, and D.~M. Tong.
\newblock Alternative framework for quantifying coherence.
\newblock {\em Physical Review A}, 94(6):060302, 2016.

\bibitem{Y18}
X.~Yuan.
\newblock Relative entropies of quantum channels with applications in resource
  theory.
\newblock arXiv preprint, 2018.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{amsthm,amsfonts,amsmath,amssymb,amscd, accents, mathrsfs, mathtools}
\usepackage{authblk}
\usepackage{hyperref}
\usepackage{multicol}
                                \usepackage{verbatim}
                                \usepackage{graphicx}

\swapnumbers
                                
\theoremstyle{plain}

\usepackage{enumitem}
\numberwithin{equation}{section}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{notation}[theorem]{Notation}
\newtheorem{example}[theorem]{Example}

\theoremstyle{remark}

\numberwithin{equation}{section}
\newcommand{\R}{{\mathbb R}}

\newcommand{\cB}{{\mathcal B}}

\newcommand{\cH}{{\mathcal H}}
\newcommand{\cL}{{\mathcal L}}
\newcommand{\cX}{{\mathcal X}}
\newcommand{\cC}{{\mathcal C}}

\newcommand{\cN}{{\mathcal N}}
\newcommand{\cM}{{\mathcal M}}
\newcommand{\cK}{{\mathcal K}}
\newcommand{\cY}{{\mathcal Y}}
\newcommand{\cI}{{\mathcal I}}

\newcommand{\cU}{{\mathcal U}}
\newcommand{\cV}{{\mathcal V}}
\newcommand{\cE}{{\mathcal E}}
\newcommand{\cD}{{\mathcal D}}

\newcommand{\ket}[1]{\left\vert #1\right\rangle}
\newcommand{\bra}[1]{\left\langle #1\right\vert}

\newcommand{\Tr}{\mathrm{Tr}}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}
\newcommand{\beann}{\begin{eqnarray*}}
\newcommand{\eeann}{\end{eqnarray*}}

\usepackage{color}

\usepackage{cite}

\begin{document}

\title{Dynamical Coherence Measures}
\author{Anna Vershynina}
\affil{\small{Department of Mathematics, Philip Guthrie Hoffman Hall, University of Houston, 
3551 Cullen Blvd., Houston, TX 77204-3008, USA}}
\renewcommand\Authands{ and }
\renewcommand\Affilfont{\itshape\small}

\date{\today}

\maketitle

\begin{abstract}
We present three measures of the dynamical coherence of channels, which are the generalization of several previous results. The measures are based on the generalized distance function between channels, which for example could be the divergence or a trace-distance. Free operations are considered to be detection-incoherent, creation-incoherent, and detection-creation incoherent. Free superchannels can be expressed using free pre- and post-processing channels. The new measures are monotone under free superchannels and convex.\end{abstract}

\section{Introduction}

Quantum coherence describes the existence of quantum interference, and it is often used in thermodynamics \cite{A14, C15, L15}, transport theory \cite{RM09, WM13}, and quantum optics \cite{G63, SZ97}, among few applications. Quantum coherence resource theory starts with free, i.e. incoherent, states, which are diagonal states in a pre-fixed basis. Free operations are some quantum channels that do not create coherence where it was absent, in other words, map the set of incoherent states to itself. Problems involving coherence included quantification of coherence \cite{BC14, RPL16, R16, SX15, V22, V23, YZ16}, distribution \cite{RPJ16}, entanglement \cite{CH16, SS15}, operational resource theory \cite{ChG16-2, CH16, DBG15, WY16},  and correlations \cite{HH18, MY16, TK16}. See \cite{SAP17} for a more detailed review.

Relatively recently, static resource theories (i.e. the one mostly concerned with states and their manipulation) have been extended to regard quantum channels as the elementary generalized resource, leading to a wide open area of research of the dynamical resource theory. Static resource theory has three main components: free states, free operations and resource measures. In analogue, dynamic resource theory must have free channels, free superchannels and resource measures. Note that the dynamical theory is a generalization of the statical one since any state can be regarded as a quantum channel mapping a trivial state to a given one. Much progress has been focused on the development of the theory of entropic quantifiers of channels and operational resource theory \cite{ChDAP08, GW21, LY20, SCG20, TEZP19, Y18}, to name a few references.

In \cite{Y18} it was shown that when the distance is measured by the relative entropy of channels (see Definition \ref{def:div}), the minimal distance to the free channels is a coherence measure of channels (see Definition \ref{def:coh-meas-dyn}). Denote $\cC$ as a set of free channels, taken to be either detection-incoherent (DI), creation-incoherent (CI), or detection-creation incoherent (DCI) (see Definition \ref{def:DCI}). Then it was shown that for a relative entropy of channels $D_{re}$, 
\begin{equation}
C^{\cC}(\cN)=\min_{\cK\in \cC}D_{re}(\cN,\cK)\ .
\end{equation}
is a coherence measure of channels. This measure is monotone for superchannels with free pre- and post-processing channels (i.e. of the form (\ref{eq:free-super-one})). 

In \cite{TEZP19} authors considered a norm $\|\cdot\|$ that is sub-multiplicative on quantum channels, sub-multiplicative with respect to the tensor products, and such that $\|\cK\|\leq 1$ for any detection-incoherent $\cK$ (e.g. a diamond norm, see Definition \ref{def:diamond-norm}), and showed that the following functional
\begin{equation}\label{eq:norm-intro}
C^{DI}(\cN)=\min_{\cK\in DI}\|\Delta\cN-\Delta\cK \| \ 
\end{equation}
is a coherence measure of channels. Here $\Delta$ is a completely dephasing operation (see Notation \ref{notation:cdo}). Additionally, it was shown that  (\ref{eq:norm-intro}) is a coherence measure when $\|\cdot\|$ is a trace-one norm (see Definition \ref{def:trace-norm}). The operational interpretation of these measures is that in a one-shot regime they appear in the probability to guess correctly if one obtained the target channel or the least distinguishable free operation, provided we can use only free measurements for the trace-one norm, and  non-free measurement are allowed on the auxiliary system for the diamond norm.

We generalize both of the above norms, and show that the following functionals are coherence measures:
\begin{align}
&C_f^{DI}(\cN)=\min_{\cK\in DI}f(\Delta\cN,\Delta\cK),\label{def:DI-f-intro}\\
&C_f^{CI}(\cN)=\min_{\cK\in CI}f(\cN\Delta,\cK\Delta),\label{def:CI-f-intro}\\
&C_f^{DCI}(\cN)=\min_{\cK\in DCI}f(\cN,\cK). \label{def:DCI-f-intro}
\end{align}
Here $f$ is a function acting as a distance between channels satisfying a particular list of properties, see Notation \ref{def:f}. As an example, this function can be taken to be a relative entropy of channels, trace-one distance, or a diamond distance. It could also be a quantum divergence 

In Chapter 2, we start with common notions and notations: relative entropy/divergence of channels, entropy of channels, static and the dynamic coherence theories. 

In Chapter 3, prove that three measures (\ref{def:DI-f-intro}), (\ref{def:CI-f-intro}), (\ref{def:DCI-f-intro}) are coherence measures of channels.

\section{Preliminaries}

In this paper, we use the following notations and definitions: all Hilbert spaces are finite dimensional. We denote them as $\cH_A, \cH_B, \dots$, where subscripts indicate the corresponding system. Sometimes we drop the subscript if the statement is true for any Hilbert space or it is evident which system we are talking about. For a Hilbert space $\cH$, we denote by $\cL(\cH)$ the space of all linear maps on $\cH$, and as $\cB(\cH)$ the space of all bounded liner operators on $\cH$.

A Hilbert space corresponding to multiple systems, e.g. AB, is a tensor product of individual subsystems, e.g. $\cH_A\otimes\cH_B$. 

For a Hilbert space $\cH_A$, its dimension is denoted as $|A|:=\dim\cH_A$. A quantum state or a density operator $\rho_A\in\cB(\cH_A)$ on a Hilbert space $\cH_A$ is a positive semidefinite, trace-normalized operator, i.e. $\rho_A\geq 0$ and $\Tr\,\rho_A=1$. A state is pure if it is rank-one. A pure state $\psi_A$ has an associated  vector $\ket{\psi}_A\in\cH_A$ such that $\langle{\psi}|\psi\rangle=1$ and $\psi_A=\ket{\psi}\bra{\psi}_A$. The set of all density operators or states on $\cH_A$ is denoted as $\cD(\cH_A)$. For a generic Hilbert space $\cH$, we denote $\cD=\cD(\cH)$.
 
A quantum channel $\cN:A\rightarrow B$ is a linear completely-positive trace-preserving (CPTP) map from $\cD(\cH_A)$ to $\cD(\cH_B)$. Sometimes the channel can also be denoted as $\cN_{A\rightarrow B}$. The identity channel on $\cD(\cH_A)$ is denoted as $I_A$. The subscript in the channels is dropped if it is evident which systems are involved. If systems $A$ and $B$ are the same, then the channel $\cN_{A\rightarrow A}$ is denoted as $\cN_A$. A composition of channels $\cN:A\rightarrow B$ and $\cK:B\rightarrow C$ is denoted as $\cK\circ\cN$, although sometimes we use notation $\cK\cN$ to simplify the formula.

A {superchannel} \cite{ChDAP08} $\Lambda$ transforms a quantum channel $\cN_{A\rightarrow B} $ to a channel from $C$ to $D$ as follows
\begin{equation}\label{def:superch}
\Lambda(\cN_{A\rightarrow B})_{C\rightarrow D}=\cM_{BE\rightarrow D}\circ(\cN_{A\rightarrow B}\otimes I_E)\circ \cK_{C\rightarrow AE}\ , 
\end{equation}
with the ancillary system $E$, and channels $\cM_{BE\rightarrow D}$ and $\cK_{C\rightarrow AE}$.

\subsection{Divergence of channels}

We consider a class of the following divergences, which is a distinguishability measure between quantum states.

\begin{definition}\label{def:div-div} A function $D:\cD\times\cD\rightarrow \R$ on a set of pairs of states is a (generalized convex) {\it divergence} if
\begin{itemize}
\item (Non-negativity) $D(\rho\|\sigma)\geq 0$, and $D(\rho\|\sigma)=0$ if and only if $\rho=\sigma$.
\item (Data processing/monotonicity). For any quantum channel $\cN$, it holds that $D(\cN(\rho)\|\cN(\sigma))\leq D(\rho\|\sigma)$.
\item (Stability) $D(\rho\otimes\tau\|\sigma\otimes\tau)=D(\rho\|\sigma)$.
\item (Joint convexity) For $0\leq \lambda\leq 1$, and quantum states $\rho_i, \sigma_i$, it holds that
$$D\left(\lambda\rho_1+(1-\lambda)\rho_2\|\lambda\sigma_1+(1-\lambda)\sigma_2\right)\leq \lambda D(\rho_1\|\sigma_1)+(1-\lambda)D(\rho_2\|\sigma_2)\ . $$ 
\end{itemize}
\end{definition}
Note that even though we explicitly state the stability property, it is a consequence of the monotonicity property, since both tensoring and the partial trace are CPTP maps (i.e. quantum channels) \cite{WWY14}. Moreover, from the monotonicity property we obtain the invariance under unitaries, i.e. $D(\rho\|\sigma)=D(U\rho U^*\|U\sigma U^*)$ for any unitary $U$.

\begin{example}
(Umegaki) Quantum relative entropy, $D_{re}(\rho\|\sigma)=\Tr(\rho\log\rho-\rho\log\sigma)$, and trace-distance, $\|\rho-\sigma\|_1=\Tr|\rho-\sigma|$, are examples of a generalized convex divergence. 
\end{example}

\begin{definition}\label{def:div}
Let $D$ be a divergence from Definition \ref{def:div-div}. Then a {\it quantum divergence of channels} $\cN_{A\rightarrow B}$ and $\mathcal{M}_{A\rightarrow B}$ based on $D$ is defined as 
\begin{equation}\label{def:qre-channel}
D(\cN \| \cM)=\sup_{\rho_{AR}} D(\cN\otimes I(\rho)\| \cM\otimes I(\rho))\ .
\end{equation}
Here the supremum is taken over all systems $R$ of any dimension and all states $\rho_{AR}$. However, it is sufficient to consider only pure states $\rho_{AR}$ with system $R$ being isomorphic to system $A$, because of the state purification, the data-processing inequality, and the Schmidt decomposition theorem. 
\end{definition}

Taking Umegaki relative entropy $D_{re}$ in the definition of the quantum divergence of channels, it becomes a relative entropy of channels, which was first proposed in \cite{CMM18}, and generalized in \cite{LKDW18}. 

When divergence is a trace-distance, then the divergence of channels is called a {\it diamond-distance} of channels, which we discuss below.

The quantum divergence of channels satisfies the following properties, which directly follows from the proofs in \cite{Y18}:
\begin{itemize}
\item (Non-negativity) $D(\cN\|\cM)\geq 0$ and $D(\cN\|\cM)=0$ if and only if $\cN=\cM$.
\item (Weak monotonicity) For any quantum channels $\mathcal{K}_i$, we have 
$$D(\mathcal{K}_1\circ\cN\circ \mathcal{K}_2\|\mathcal{K}_1\circ\cM\circ \mathcal{K}_2 )\leq D(\cN\|\cM)\ .$$
\item (Strong monotonicity) For any super-channel $\Lambda$, 
$$D(\Lambda(\cN)\|\Lambda(\cM))\leq D(\cN\|\cM)\ .$$
\item (Joint convexity) For $0\leq\lambda\leq 1$, and quantum channels $\cN_i, \cM_i$, we have $$D\left(\lambda\cN_1+(1-\lambda)\cN_2\|\lambda\cM_1+(1-\lambda)\cM_2\right)\leq \lambda D(\cN_1\|\cM_1)+(1-\lambda)D(\cN_2\|\cM_2)\ .$$
%\item (Additivity) $D(\cN_1\otimes\cN_2\|\cM_1\otimes\cM_2)\geq D(\cN_1\|\cM_1)+D(\cN_2\|\cM_2)$.
\item (Stability) $D( \cN\otimes I\| \cM\otimes I)=D(\cN\|\cM)$.
 \end{itemize}
 The properties of the quantum divergence of channels is discussed in Chapter 3. 

\subsection{Trace and diamond norms}

\begin{definition}\label{def:trace-norm} \cite{Wat18} The {\it trace-norm of a linear map} $\cX:\cL(\cH_A)\rightarrow\cL(\cH_B)$ is defined as 
\begin{equation}\label{def:trace-norm-channel}
\|\cX\|_1=\max \{ \|\cX(X)\|_1\ :\ X\in\cL(\cH_A),\ \|X\|_1\leq 1\}  \ ,
\end{equation}
where the trace-norm on $\cL(\cH)$ is defined as $\|M\|_1=\Tr\sqrt{M^*M}$ for $M\in\cL(\cH)$. Note that the maximum above is always achieved by a rank-one operator (Proposition 3.38 \cite{Wat18}). For a positive map the maximum is achieved by a rank-one projection.
\end{definition}
\begin{definition}
The {\it trace-distance of quantum channels} $\cN, \cM:A \rightarrow B$ is defined as 
\begin{equation}\label{def:trace-norm-channel}
\|\cN -\cM\|_1=\max_{\rho\in\cD(\cH_A)} \|\cN(\rho) - \cM (\rho)\|_1\ .
\end{equation}
\end{definition}

The trace-norm of a quantum channel has the following properties, (for more properties see \cite{Wat18}):
\begin{enumerate}
\item The trace-norm of a quantum channel $\cN$ is one, i.e. $\|\cN\|_1=1$.
\item For all linear maps $\cX: \cL(\cH_A)\rightarrow\cL(\cH_B)$, $\cY: \cL(\cH_B)\rightarrow\cL(\cH_C)$, it holds $\|\cY\cX\|_1\leq \|\cY\|_1\|\cX\|_1$. 
\item For all channels $\cN_1,\cM_1:A\rightarrow B$ and $\cN_2, \cM_2:B\rightarrow C$, it holds that
$$\|\cN_2\cN_1-\cM_2\cM_1 \|_1\leq \|\cN_1-\cM_1\|_1+\|\cN_2-\cM_2\|_1\ .$$
\end{enumerate}

\begin{definition}\label{def:diamond-norm} \cite{Wat18}
The {\it diamond norm (or the completely bounded trace-norm) of a linear map} $\Omega: \cL(\cH_A)\rightarrow\cL(\cH_B)$ is defined as 
$$\|\Omega_{A\rightarrow B}\|_\diamond=\|\Omega_{A\rightarrow B}\otimes I_R\|_1\ , $$
where $\dim R=\dim A$.
%It is sufficient to consider only systems $R$ isomorphic to the system $A$ .  
\end{definition}

The diamond norm satisfies the following properties in particular, \cite{Wat18}:
\begin{enumerate}
\item The diamond norm of a quantum channel $\cN$ is one, i.e. $\|\cN\|_\diamond=1$. 
\item Since trace-norm is monotone under quantum channels, in particular partial traces, we get that for any linear map $\Omega$, 
$$ \|\Omega\|_1\leq \|\Omega\|_\diamond\ . $$
\item (Sub-multiplicativity) for any linear maps $\Omega: \cL(\cH_B)\rightarrow\cL(\cH_C)$ and $\Sigma: \cL(\cH_A)\rightarrow\cL(\cH_B)$
$$\|\Omega\Sigma\|_\diamond\leq \|\Omega\|_\diamond\|\Sigma\|_\diamond\ . $$
\item For all channels $\cN_1,\cM_1:A\rightarrow B$ and $\cN_2, \cM_2:B\rightarrow C$, it holds that
$$\|\cN_2\cN_1-\cM_2\cM_1 \|_\diamond\leq \|\cN_1-\cM_1\|_\diamond+\|\cN_2-\cM_2\|_\diamond\ .$$
\item (Multiplicativity under tensor products) for any linear maps $\Omega: \cL(\cH_A)\rightarrow\cL(\cH_B)$ and $\Sigma: \cL(\cH_C)\rightarrow\cL(\cH_D)$
$$\|\Omega\otimes\Sigma\|_\diamond= \|\Omega\|_\diamond\|\Sigma\|_\diamond\ . $$
\item (Monotonicity under superchannels)  Let $\Lambda$ be a superchannel (\ref{def:superch}). Then from the above propertires it follows that
$$\|\Lambda(\cN)\|_\diamond\leq \|\cM\|_\diamond \|\cN\otimes I\|_\diamond\|\cK\|_\diamond=\|\cN\|_\diamond \ ,$$
since $\cM$ and $\cK$ are channels,  their diamond norm is one, and since $\|\cN\otimes I\|_\diamond=\|\cN\|_\diamond$.
\end{enumerate}

\begin{definition}
Taking the diamond-distance as a divergence in (\ref{def:div}) defines a {\it diamond-distance of channels},
\begin{equation}\label{def:qre-channel}
\|\cN - \cM\|_\diamond=\max_{\rho} \|\cN\otimes I(\rho)- \cM\otimes I(\rho)\|_1\ .
\end{equation}
\end{definition}

Recall that for the Umegaki relative entropy the Pinsker's inequality for states holds: for any  states $\rho, \sigma$, we have
$$D(\rho\|\sigma)\geq \frac{1}{2}\|\rho-\sigma\|_1^2\ . $$

Straight from definition of the quantum relative entropy of channels,  (\ref{def:qre-channel}), and the Pinsker's inequality, the Pinsker's inequality for channels holds.
\begin{proposition}
For any quantum channels $\cN, \cM$, the Pinsker's inequality for channels holds
$$D_{re}(\cN\|\cM)\geq \frac{1}{2}\|\cN-\cM\|_\diamond^2\geq \frac{1}{2}\|\cN-\cM\|_1^2\ . $$
Here $D_{re}$ is the relative entropy of channels based on the Umegaki relative entropy.
\end{proposition}

\subsection{Static resource theory of coherence}

Let $\cH$ be a $d$-dimensional Hilbert space. Let us fix an orthonormal basis $\cE=\{\ket{j}\}_{j=1}^d$ of vectors in $\cH$.
\begin{definition} A state $\delta$ is called {\it incoherent} in $\cE$ if it can be represented as follows
$\delta=\sum_j \delta_j\ket{j}\bra{j}. $ Denote the set of {\bf incoherent states} for a fixed basis $\cE=\{\ket{j}\}_j$ as $\cI_\cE=\{\rho=\sum_jp_j\ket{j}\bra{j}\}.$ We drop the subscript $\cE$ from now on.
\end{definition}

\begin{notation}\label{notation:cdo}
The {\it completely dephasing operator}  $\Delta_\cE: \mathcal{D}(\cH)\rightarrow \cI$ is defined as
\begin{equation}\label{def:dephasing}
\Delta_\cE(\rho)=\sum_j \bra{j}\rho\ket{j} \ket{j}\bra{j}\ ,
\end{equation}
Sometimes we drop the subscript $\cE$, or use a subscript denoting the system this operator acts on.
\end{notation}

Any quantum channel $\Phi$ admits a Kraus decomposition (which may not be unique) of the form
\begin{equation}\label{eq:Kraus}
\cN(\rho)=\sum_n K_n \rho K_n^*\ , \ \ \sum_n K_n^*K_n=I\ .
\end{equation}

\begin{definition} 
A quantum channel $\cN$ on $\cD(\cH)$ is an {\it incoherent operation (IO)}, if there exists a Kraus decomposition (\ref{eq:Kraus}) such that  $K_n \cI K_n^*\subset \cI,\ \text{for all }n.$
\iffalse
\begin{itemize}
\item a {\it maximal incoherent operation (MIO)}, if $\cN( \cI) \subset \cI.$
\item an {\it incoherent operation (IO)}, if there exists a Kraus decomposition (\ref{eq:Kraus}) such that  $K_n \cI K_n^*\subset \cI,\ \text{for all }n.$
\item a {\it dephasing-covariant incoherent operation (DIO)}, if $\cN\Delta=\Delta\cN.$
\item  a {\it strictly incoherent operation (SIO)}, if if there exists a Kraus decomposition
(\ref{eq:Kraus})  such that
$K_n \Delta(\rho) K_n^*=\Delta\left(K_n\rho K_n^* \right),\ \text{for all }n.$ 
\end{itemize}
\fi
\end{definition}

\begin{definition}
A {\it measure of coherence} $\cC(\rho)$ should satisfy the following conditions
\begin{itemize}
\item (C1) $\cC(\rho)\geq 0$, and $\cC(\rho)=0$ if and only if $\rho\in\cI$;
\item (C2) Non-selective monotonicity under IO (monotonicity): for all IO $\cN$ and all states $\rho$,
$$\cC(\rho)\geq \cC(\cN(\rho))\ ; $$
\item (C3) Selective monotonicity under IO (strong monotonicity): for all IO $\cN$ with Kraus operators $K_n$, and all states $\rho$,
$$\cC(\rho)\geq \sum_n p_n \cC(\rho_n)\ , $$
where $p_n$ and $\rho_n$ are the outcomes and post-measurement states
$$\rho_n=\frac{K_n\rho K_n^*}{p_n},\ \ p_n=\Tr K_n\rho K_n^*\ . $$
\item (C4) Convexity, 
$$\sum_n p_n \cC(\rho_n)\geq \cC\left(\sum_n p_n\rho_n\right)\ , $$
for any sets of states $\{\rho_n\}$ and any probability distribution $\{p_n\}$.
\end{itemize}
\end{definition}
These properties are parallel with the entanglement measure theory, where the average entanglement is not increased under the local operations and classical communication (LOCC). Notice that coherence measures that satisfy conditions (C3) and (C4) also satisfies condition (C2) \cite{BC14}. 

While properties (C2) and (C3) are formulated for incoherent operations, one may consider different classes of incoherent operations (such as maximally incoherent operations, strictly incoherent operations, and many others, see for example  \cite{ChG16, ChG16-2, DevS16} for their comparison and analysis).

\subsection{Dynamic resource theory of coherence}

Quantum channels now play the role of states in the static resource theory. While there is only one commonly agreed set of free incoherent states (diagonal in the pre-fixed basis), there are multiple way one can consider free channels in the dynamic resource theory, such as classical channels, or detection/creation/detection-creation incoherent channels \cite{SCG20, TEZP19}. 

Let $\cN:A\rightarrow B$ be a quantum channel, and let us fix bases $\cE_A$ and $\cE_B$ in systems $A$ and $B$ respectively. Then $\Delta_{\cE_A}$ and $\Delta_{\cE_B}$ denote the completely dephasing operators (\ref{def:dephasing}) in systems $A$ and $B$ resp.

\begin{definition}\label{def:DCI}\cite{SCG20, TEZP19}
A quantum channel $\cN_{A\rightarrow B}$ is
\begin{itemize}
\item {\it detection incoherent (DI)} iff $\Delta_{\cE_B}\cN_{A\rightarrow B}=\Delta_{\cE_B}\cN_{A\rightarrow B}\Delta_{\cE_A},$ 
\item {\it creation incoherent (CI)} iff $\cN_{A\rightarrow B}\Delta_{\cE_A}=\Delta_{\cE_B}\cN_{A\rightarrow B}\Delta_{\cE_A},$
\item {\it detection-creation incoherent (DCI) or DIO} iff $\Delta_{\cE_B}\cN_{A\rightarrow B}=\cN_{A\rightarrow B}\Delta_{\cE_A}$.
\end{itemize}
See \cite{TEZP19} for more detailed information on the structure of these channels.
\end{definition}

A set of free superchannels can be defined in multiple ways \cite{SCG20, TEZP19, Y18}. 
Similar to \cite{TEZP19, Y18}, we consider the following free superchannels. 

\begin{notation}
A superchannel $\Lambda$ is free if there is a decomposition 
\begin{equation}\label{eq:free-super-one}
\Lambda(\cN_{A\rightarrow B})=\cK_{BE\rightarrow D}\circ(\cN_{A\rightarrow B}\otimes I_{E})\circ\cM_{C\rightarrow AE}\ ,
\end{equation}
where $\cM_{C\rightarrow AE}, \cK_{BE\rightarrow D}$ are free channels taken to be either DI, CI, or DCI.
\end{notation}

\begin{definition}\label{def:coh-meas-dyn} \cite{TEZP19, Y18}
{\it Coherence measures} are real-valued functions $C$ of channels $\cN$ satisfying the following conditions:
\begin{itemize}
\item (C1) {Non-negativity:} $C(\cN)\geq 0$ and $C(\cN)=0$ for free channels $\cN$.
\item (C2) {Monotonicity:} for free superchannels $\Lambda$, the coherence measure is non-increasing, $C(\Lambda(\cN))\leq C(\cN)$.
\end{itemize}
Sometimes other conditions are included as well, for example:
\begin{itemize}
\item (C3) Convexity: for channels $\cN, \cK$ and $0\leq \lambda\leq 1$, it holds $C(\lambda \cN+(1-\lambda)\cK)\leq \lambda C(\cN)+(1-\lambda) C(\cK).$
\end{itemize}
\end{definition}

\section{Coherence measures of channels}

\begin{notation}\label{def:f}
Let function $f(\cN,\cM)$ between two channels $\cN, \cM$, viewed as a distance between them, be such that 
\begin{itemize}
\item (Non-negativity) For quantum channels $\cN, \cM$,  it holds that $f(\cN,\cM)\geq 0$, and $f(\cN,\cM)=0$ if and only if $\cN=\cM$.
\item (Weak monotonicity) For quantum channels $\cV,\cU,\cN, \cM$, the function is non-increasing under composition $f(\cV\circ\cN\circ\cU, \cV\circ\cM\circ\cU)\leq f(\cN, \cM)$. %It is possible to restrict this condition for when $\cV, \cU$ are free (DI, CI, or DCI) channels only.
\item (Joint convexity). For $0\leq \lambda\leq 1$, it holds that $f\left(\lambda\cN_1+(1-\lambda)\cN_2, \lambda\cM_1+(1-\lambda)\cM_2\right)\leq \lambda f(\cN_1, \cM_1)+(1-\lambda)f(\cN_2,\cM_2)$.
\item (Monotone under tensor product with the identity operator). For quantum channels $\cN, \cM$,  it holds that $f(\cN\otimes I,\cM\otimes I)\leq f(\cN, \cM)$.
\item (Monotone under tensor product with the dephasing operator). For quantum channels $\cN, \cM$, and the completely dephasing operator $\Delta$, it holds that $f(\cN\otimes\Delta,\cM\otimes \Delta)\leq f(\cN, \cM)$.
\end{itemize}
\end{notation}

The following lemma can be obtained from the proof of Theorem 9 in \cite{TEZP19}.
\begin{lemma}
Trace-one distance of channels, $f(\cN,\cM)=\|\cN-\cM\|_1$, satisfy the above conditions.
\end{lemma}

The proof of the following lemma is similar to the proof of the lemma above, but we provide it for the completeness sake below.
\begin{lemma}\label{lemma:RE}
Quantum divergence of channels (see Definition \ref{def:div}), $f(\cN,\cM)=D(\cN\|\cM)$, satisfy the above conditions for any divergence from Definition \ref{def:div-div}. In particular, the relative entropy of channels and the diamond-distance of channels are such functions.
\end{lemma}
\begin{proof}
We only need to check the monotonicity under tensor product with the dephasing operator, since other properties are satisfied by \cite{Y18}.

Let $\cN_{A\rightarrow B}, \cM_{A\rightarrow B}$ be two quantum channels. For a system $C$ with a fixed basis $\cE_C$, denote the completely dephasing operator (\ref{def:dephasing}) as $\Delta_C=\Delta_{\cE_C}$. Then
\begin{align}
D(\cN_{A\rightarrow B}\otimes\Delta_C\|\cM_{A\rightarrow B}\otimes\Delta_C)&=\max_{\rho_{ACR}}D(\cN_{A\rightarrow B}\otimes\Delta_C\otimes I_R(\rho_{ACR})\|\cM_{A\rightarrow B}\otimes\Delta_C\otimes I_R(\rho_{ACR}))\\
&=\max_\rho D([\cN\otimes I\otimes I][I\otimes \Delta\otimes I ](\rho)\|[\cM\otimes I\otimes I][I\otimes \Delta\otimes I] (\rho))\\
&=\max_{\sigma_{ACR}=I_A\otimes \Delta_C\otimes I_R(\rho)} D(\cN\otimes I\otimes I(\sigma)\|\cM\otimes I\otimes I(\sigma))\ .
\end{align}

Any state $\rho_{ACR}$ can be written as $\rho_{ACR}=\sum_{i,j,a,b}\alpha_{i,j,a,b}\ket{a}\bra{b}_{AR}\otimes\ket{i}\bra{j}_C$, where $\{\ket{i}\}$  (or $\{\ket{j}\}$) is the fixed basis of the operator $\Delta$. Then
\begin{align}
I_A\otimes \Delta_C\otimes I_R (\rho_{ACR})&=\sum_{i,a,b}\alpha_{i,i,a,b}\ket{a}\bra{b}_{AR}\otimes\ket{i}\bra{i}_C\\
&=:\sum_i p_i \, \sigma_i\otimes \ket{i}\bra{i}\ ,
\end{align}
here we denoted  $p_i\sigma_i:=\sum_{a,b}\alpha_{i,i,a,b}\ket{a}\bra{b}$ such that $\Tr\sigma_i=1$ and $\sum_ip_i=1$.

Therefore,
\begin{align}
D(\cN\otimes \Delta\|\cM\otimes \Delta)&=\max_{\sigma_{i_{AR}}, p_i} D\left(\sum_i p_i [\cN\otimes I](\sigma_i)_{AR}\otimes \ket{i}\bra{i}_C\|\sum_ip_i[\cM\otimes I](\sigma_i)_{AR}\otimes \ket{i}\bra{i}_C\right)\\
&\leq \max_{\sigma_{i_{AR}}, p_i} \sum_ip_i D([\cN\otimes I](\sigma_i)_{AR}\otimes \ket{i}\bra{i}_C\|[\cM\otimes I](\sigma_i)_{AR}\otimes \ket{i}\bra{i}_C)\\
&=\max_{\sigma_i, p_i} \sum_ip_i D(\cN\otimes I(\sigma_i)\|\cM\otimes I(\sigma_i))\\
&\leq \max_{\sigma_{AR}} D(\cN_{A\rightarrow B}\otimes I_R(\sigma_{AR})\|\cM_{A\rightarrow B}\otimes I_R(\sigma_{AR}))\\
&=D(\cN\|\cM)\ .
\end{align}

In the first inequality we used joint convexity of quantum relative entropy. In the second equality we used stability of the divergence.
\end{proof}

\begin{theorem}
Let $f$ be a function satisfying the properties in Notation \ref{def:f}. Consider three cases, when sets DI, CI, and DCI, (see Definition \ref{def:DCI}), are considered to be sets of free operations. Define:
\begin{align}
&C_f^{DI}(\cN)=\min_{\cK\in DI}f(\Delta\cN,\Delta\cK),\label{def:DI-f}\\
&C_f^{CI}(\cN)=\min_{\cK\in CI}f(\cN\Delta,\cK\Delta),\label{def:CI-f}\\
&C_f^{DCI}(\cN)=\min_{\cK\in DCI}f(\cN,\cK). \label{def:DCI-f}
\end{align}

These functionals are coherence measures of channels, i.e. they satisfy three conditions in Definition \ref{def:coh-meas-dyn}. The free superchannels are  of the type (\ref{eq:free-super-one}) with free channels taken from a corresponding set of free operations (DI, CI, or DCI).

%\begin{enumerate}[label=\Roman*]
%\item $C_f^\cC(\cN)=0$ if and only if $\cN\in \cC$.
%\item Let $\Lambda$ be a free superchannel of the either type (\ref{eq:free-super-one}) or (\ref{eq:free-super-two}) with free channels taken from $\cC$. Then the coherence measure is monotone under $\Lambda$,
%$$C_f^\cC(\Lambda(\cN))\leq C_f^\cC(\cN)\ .$$
%\item For any quantum channels  $\cN, \cK$  and  $0\leq\lambda\leq 1$, the coherence measure is convex
%$$C_f^\cC(\lambda \cN+(1-\lambda)\cK)\leq \lambda C_f^\cC(\cN)+(1-\lambda)C_f^\cC(\cK)\ . $$
%\end{enumerate}
\end{theorem}

Note that one may think that all three coherence measures (\ref{def:DI-f})-(\ref{def:DCI-f}) can be viewed as the minimal distance to a free set, e.g. one can define $C(\cN)=\min_{\cK\in DI}F(\cN,\cK)$, with the distance function defined as 
$F(\cN,\cM)=f(\Delta\cN,\Delta\cK)$, and similarly for CI. However, the simplest property, $F(\cN,\cM)=0$ if and only if $\cN=\cM$, is not true. Therefore (\ref{def:DI-f}) and (\ref{def:CI-f}) are not distance measures.

\begin{proof} Denote the set of free operations DI, CI, or DCI as $\cC$.

1. If $\cN\in\cC$, then $C_f^\cC(\cN)=0$. On the other hand, if $C_f^\cC(\cN)=0$, then there exists $\cK\in\cC$ such that
\begin{itemize}
\item for $\cC=DI$: $\Delta\cN=\Delta\cK=\Delta\cK\Delta=\Delta\cN\Delta$, therefore $\cN\in DI$.
\item for $\cC=CI$: $\cN\Delta=\cK\Delta=\Delta\cK\Delta=\Delta\cN\Delta$, therefore $\cN\in CI$.
\item for $\cC=DCI$: $\cN=\cK\in DCI$.
\end{itemize}

2. {Note that a composition of two maps in $\cC$ is in $\cC$.} I.e. if $\cK_1, \cK_2\in DI$, then $\cK_1\cK_2\in DI$ since
\begin{align}
\Delta\cK_1\cK_2&=\Delta\cK_1\Delta\cK_2\\
&=\Delta\cK_1\Delta\cK_2\Delta\\
&=\Delta\cK_1\cK_2\Delta\ .
\end{align}
The first equality holds since $\cK_1\in DI$, i.e. $\Delta\cK_1=\Delta\cK_1\Delta$. The second equality holds since $\cK_2\in DI$, i.e. $\Delta\cK_2=\Delta\cK_2\Delta$. The last equality holds since again $\cK_1\in DI$ and therefore $\Delta\cK_1\Delta=\Delta\cK_1$.

Similarly, if $\cK_1, \cK_2\in CI$, then $\cK_1\cK_2\in CI$ since
\begin{align}
\cK_1\cK_2\Delta&=\cK_1\Delta\cK_2\Delta\\
&=\Delta\cK_1\Delta\cK_2\Delta\\
&=\Delta\cK_1\cK_2\Delta\ .
\end{align}
The first equality holds since $\cK_2\in CI$, i.e. $\cK_2\Delta=\Delta\cK_2\Delta$. The second equality holds since $\cK_1\in CI$, i.e. $\cK_1\Delta=\Delta\cK_1\Delta$. The last equality holds since again $\cK_2\in CI$ and therefore $\Delta\cK_2\Delta=\cK_2\Delta$.

Similarly, if $\cK_1, \cK_2\in DI$, then $\cK_1\cK_2\in DCI$ since
\begin{align}
\Delta\cK_1\cK_2&=\cK_1\Delta\cK_2\\
&=\cK_1\cK_2\Delta\ .
\end{align}
The first equality holds since $\cK_1\in DCI$, i.e. $\Delta\cK_1=\cK_1\Delta$. The second equality holds since $\cK_2\in DI$, i.e. $\Delta\cK_2=\cK_2\Delta$.

3. Let $\Phi\in DI$ be any channel in DI. Then
\begin{align}
C_f^{DI}(\cN\circ \Phi)&=\min_{\cK\in DI}f(\Delta\cN\Phi,\Delta\cK)\\
&\leq\min_{\cK'\in DI}f(\Delta\cN\Phi,\Delta\cK'\Phi)\\
&\leq\min_{\cK'\in DI}f(\Delta\cN,\Delta\cK')\\
&= C_f^{DI}(\cN)\ .
\end{align}
In the first inequality we used that the composition of DI maps is DI. In the second inequality we used that the function $f$ is weakly monotone for quantum channels.

Let $\Phi\in CI$ be any channel in CI. Then
\begin{align}
C_f^{CI}(\cN\circ \Phi)&=\min_{\cK\in CI}f(\cN\Phi\Delta,\cK\Delta)\\
&\leq\min_{\cK'\in CI}f(\cN\Phi\Delta,\cK'\Phi\Delta)\\
&=\min_{\cK'\in CI}f(\cN\Delta\Phi\Delta,\cK'\Delta\Phi\Delta)\\
&\leq\min_{\cK'\in CI}f(\cN\Delta,\cK'\Delta)\\
&= C_f^{CI}(\cN)\ .
\end{align}
In the first inequality we used that the composition of CI maps is CI. The last inequality is true, since the function $f$ is weakly monotone on quantum channels.

Let $\Phi\in DCI$ be any channel in DCI. Then
\begin{align}
C_f^{DCI}(\cN\circ \Phi)&=\min_{\cK\in DCI}f(\cN\Phi,\cK)\\
&\leq\min_{\cK'\in DCI}f(\cN\Phi,\cK'\Phi)\\
&\leq\min_{\cK'\in DCI}f(\cN,\cK')\\
&= C_f^{DCI}(\cN)\ .
\end{align}
In the first inequality we used that the composition of DCI maps is DCI. The last inequality is true, since the function $f$ is weakly monotone on quantum channels.

4.  Let $\Phi\in DI$. Then
\begin{align}
C_f^{DI}(\Phi\circ \cN)&=\min_{\cK\in DI}f(\Delta\Phi\cN,\Delta\cK)\\
&\leq \min_{\cK'\in DI}f(\Delta\Phi\cN,\Delta\Phi\cK')\\
&= \min_{\cK'\in DI}f(\Delta\Phi\Delta\cN,\Delta\Phi\Delta\cK')\\
&\leq \min_{\cK'\in DI}f(\Delta\cN,\Delta\cK')\\
&=C_f^{DI}(\cN)\ .
\end{align} 
The last inequality is true, since $f$ is weakly monotone.

Let $\Phi\in CI$. Then
\begin{align}
C_f^{CI}(\Phi\circ \cN)&=\min_{\cK\in CI}f(\Phi\cN\Delta,\cK\Delta)\\
&\leq \min_{\cK'\in CI}f(\Phi\cN\Delta,\Phi\cK'\Delta)\\
&\leq \min_{\cK'\in CI}f(\cN{\Delta},\cK'{\Delta})\\
&=C_f^{CI}(\cN)\ .
\end{align} 
The last inequality is true, since $f$  is weakly monotone.

Let $\Phi\in DCI$. Then
\begin{align}
C_f^{DCI}(\Phi\circ \cN)&=\min_{\cK\in DCI}f(\Phi\cN,\cK)\\
&\leq \min_{\cK'\in DCI}f(\Phi\cN,\Phi\cK')\\
&\leq \min_{\cK'\in DCI}f(\cN,\cK')\\
&=C_f^{DCI}(\cN)\ .
\end{align} 
The last inequality is true, since $f$ is weakly monotone.

5. Consider
\begin{align}
C_f^{DI}(\cN\otimes I)&=\min_{\cK\in DI}f(\Delta(\cN\otimes I),\Delta\cK)\\
&\leq \min_{\cK=\cK'\otimes I\in DI}f(\Delta(\cN\otimes I),\Delta(\cK'\otimes I))\\
&=\min_{\cK'\in DI}f(\Delta\cN\otimes \Delta,\Delta\cK'\otimes \Delta)\\
&\leq \min_{\cK'\in DI}f(\Delta\cN,\Delta\cK')\\
&=C_f^{DI}(\cN)\ .
\end{align}
The last inequality holds since $f$ is monotone under tensor product with the dephasing operator.

Similarly,
\begin{align}
C_f^{CI}(\cN\otimes I)&=\min_{\cK\in CI}f((\cN\otimes I)\Delta,\cK\Delta)\\
&\leq \min_{\cK=\cK'\otimes I\in CI}f((\cN\otimes I)\Delta,(\cK'\otimes I)\Delta)\\
&=\min_{\cK'\in CI}f(\cN\Delta\otimes \Delta,\cK'\Delta\otimes \Delta)\\
&\leq \min_{\cK'\in CI}f(\cN\Delta,\cK'\Delta)\\
&=C_f^{CI}(\cN)\ .
\end{align}

Also,
\begin{align}
C_f^{DCI}(\cN\otimes I)&=\min_{\cK\in DCI}f(\cN\otimes I,\cK)\\
&\leq \min_{\cK=\cK'\otimes I\in DCI}f(\cN\otimes I,\cK'\otimes I)\\
&\leq \min_{\cK'\in DCI}f(\cN,\cK')\\
&=C_f^{DCI}(\cN)\ .
\end{align}
In the last inequality we used that $f$ is monotone under tensor product with the identity operator.

From 2-5, for any free quantum superchannel $\Lambda(\cN_{A\rightarrow B})=\Phi_{BE\rightarrow D}\circ(\cN_{A\rightarrow B}\otimes I_{E})\circ\Psi_{C\rightarrow AE}$ with free $\Phi, \Psi\in \cC$, all coherence measures are monotone,
$$C_f^\cC(\Lambda(\cN))\leq C_f^\cC(\cN)\ . $$

6. Let $0\leq \lambda\leq 1$. Denote $\cK_1\in DI$ and $\cK_2\in DI$ as channels such that
$$C_f^{DI}(\cN)=f(\Delta\cN, \Delta\cK_1)\ ,\ \ \  C_f^{DI}(\cK)=f(\Delta\cK,\Delta\cK_2)\ .$$ 
Then
\begin{align}
C_f^{DI}(\lambda\cN+(1-\lambda)\cK )&=\min_{\Phi\in DI}f\left(\Delta(\lambda\cN+(1-\lambda)\cK ),\Delta\Phi\right)\\
&\leq f\left(\Delta(\lambda\cN+(1-\lambda)\cK ),\Delta(\lambda\cK_1+(1-\lambda)\cK_2)\right)\\
&=f\left(\lambda \Delta\cN+(1-\lambda)\Delta\cK ,\lambda\Delta\cK_1+(1-\lambda)\Delta\cK_2\right)\\
&\leq \lambda f(\Delta\cN, \Delta\cK_1)+(1-\lambda)f(\Delta\cK,\Delta\cK_2)\\
&=\lambda C_f^{DI}(\cN)+(1-\lambda)C_f^{DI}(\cK)\ .
\end{align}
Here, in the second inequality we used the joint convexity of $f$.

{Similar results are straightforward for CI and DCI.}
\end{proof}

\vspace{0.3in}
\textbf{Acknowledgments.}  A. V. is supported by NSF grant DMS-2105583.
\vspace{0.3in}

\bibliography{Bibliography-Vershynina}{}
\bibliographystyle{plain}

\end{document}